# Sample-Efficiency in Multi-Batch Reinforcement Learning: The Need for Dimension-Dependent Adaptivity

### v1

@@ 2 Preliminaries
A Markov Decision Process (MDP) (Puterman, 1994) is a discrete-time stochastic process comprised of a set of states $\mathcal{S}$, a set of actions $\mathcal{A}=\bigcup_{s\in\mathcal{S}}\{\mathcal{A}_{s}\}$ where $\mathcal{A}_{s}$ is the action space in state $s\in\mathcal{S}$ and, for each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}_{s}$, a next-state transition function given by a measure $p(\cdot|s,a)$ and a (deterministic) reward function $r(s,a)\in[-1,1]$. In a state $s$, an agent chooses an action $a$, receives a reward $r(s,a)$ and transitions to a new state according to $p(\cdot|s,a)$. Once in the new state, the process continues. The actions chosen by an agent are formalised by policies. A deterministic policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ is a mapping from a state to an action. In each state $s\in\mathcal{S}$, an agent following policy $\pi$ chooses action $\pi(s)\in\mathcal{A}_{s}$. We do not consider stochastic policies.
@@ 4 Main Results
In this section, we present our main results: lower-bounds on the number of rounds $K$ for sample-efficient algorithms. First, we state some assumptions. We assume $\gamma>\sqrt{3/4}$.
We consider a linear representation of the action-values for a known feature map of state-action pairs. This is a form of linear function approximation that is strictly more general than linear MDPs (Zanette et al., 2020), which assume the reward and transition functions are linearly representable. Specifically, we consider the following assumptions:
@@
This first assumption is only used for results on PE problems.
@@
The second assumption is stronger as the action-value of every policy (not just policies in $\Pi$) is linearly represented and in particular it holds for $Q^{\star}$ as it corresponds to the action-value of the optimal policy $\pi^{\star}$. We assume that when these assumptions hold, the learner is aware of them.
Finally, we formally define a sample-efficient learner:
@@ 4.1 Policy-Induced Queries
We first present a result for PE under policy-induced queries. Since all MDPs in the class used in the proof share the same dynamics, policy-induced queries are equivalent to policy-free queries (see Remark 3.6) and the result holds for both. The full proof can be found in Appendix D.
@@
In the class of MDPs used for Theorem 4.4 we can hide information about $M\in\mathcal{M}$ in the target-policy for PE but cannot for BPI. Instead, we could hide information in the transitions but this can be revealed by following policy trajectories (policy-induced queries) in our constructions. In the next section, we restrict the learner to policy-free queries and provide results for both PE and BPI.
@@ 4.2 Policy-Free Queries
We now consider only policy-free queries, which gives the environment freedom to hide information in the transition function of the MDP and leads to lower-bounds for PE and BPI that hold for the stronger Assumption 4.2 of all-policy realizability. The full proof can be found in Appendix E.
@@ 4.3 Discussion
Furthermore, in Appendix A we provide results for the fully adaptive setting where the number of batches $K$ is equal to the number of total queries $n$ (i.e. each batch contains a single query). We show that if the feature-space covers $\mathcal{B}$, there is a learner that solves any realizable (Assumption 4.1) PE problem in $d$ queries. Assuming a known target policy $\pi$, a linear dependence on $d$ was already known to be possible using roll-outs from $\pi$ (Lattimore et al., 2020), however the dependence on $d$ was coupled with other quantities such as the effective horizon $1/(1-\gamma)$ and desired accuracy $\varepsilon$. Our result states that $d$-queries are sufficient to find $Q^{\pi}$ exactly ($\varepsilon=0$), independently of $\gamma$. Our result relies heavily on the condition that the learner observes the transition function $p_{M}(\cdot|s,a)$ rather than a sample $s^{\prime}\sim p_{M}(\cdot|s,a)$ for a query $(s,a)$ (see Section 3.3), though our result under this condition is strong since using this condition with the roll-outs from $\pi$ would not give exact convergence in $d$-queries. We also provide a matching lower-bound (that also holds for BPI). These results serve to illustrate the trade-off between sample-efficiency and low-adaptivity for PE under our framework. Full-adaptivity allows low sample-complexity, while reducing adaptivity below $K=o(\log\log d)$ comes at the cost of high sample-complexity (losing sample-efficiency).
@@ 5 Proof Sketch
In this section, we provide intuition for the proof of Theorem 4.4 regarding the hardness of learning under the framework described in Section 3. We extend the ideas of Zanette (2021) beyond offline RL to our multi-batch problem. We consider the PE problem with policy-free queries under Assumption 4.1 for feature vectors $\phi(\cdot,\cdot)$ covering the unit Euclidean ball $\mathcal{B}$ (see Figure 1). The intuition for Theorem 4.5 is closely related.
@@
where $\mathcal{C}_{\gamma}(w)=\{x\in\mathcal{B}:x^{T}w/\|w\|_{2}>\gamma\}$ is the $\gamma$-hyperspherical cap of $w$. For a $w\in\partial\mathcal{B}$ and a carefully designed target policy $\pi_{w}$, we show that Assumption 4.1 holds. We also show with some of the reasoning described earlier in this section that if $n_{k}=poly(d)$ for all $k$ and $K=o(\log\log d)$, then none of the learner’s queries are in $\mathcal{C}_{\gamma}(w)\cup\mathcal{C}_{\gamma}(-w)$. Therefore the feedback observed contains only rewards of 00 and no information about the sign of the rewards in $\mathcal{C}_{\gamma}(w)\cup\mathcal{C}_{\gamma}(-w)$. $M_{w,+}$ is indistinguishable from $M_{w,-}$. Since $Q_{M_{w,+}}^{\pi_{w}}(\bar{s},w)=1$ and $Q_{M_{w,-}}^{\pi_{w}}(\bar{s},w)=-1$, the learner must incur an error of $1$ with probability at least $1/2$. We provide all the details in Appendix D. The construction for Theorem 4.5 is similar but the transitions differ across MDPs (see Appendix E).
6 Related Works
@@ 7 Conclusion
It remains unclear if the $\log\log d$ dependence on $d$ is tight. An upper-bound similar to the one we have given for the fully-adaptive PE problem in Appendix A could be established by developing new tools in the theory of subspace covering. These would formalise the number of directions for which the learner can prevent information erasure. It is also unclear if Theorem 4.4 under policy-induced queries also holds for BPI or if the sample-efficiency of BPI-algorithms in low-adaptivity settings differs for policy-induced and policy-free queries. We leave these as future work.
Appendix A Bounds for the Fully Adaptive Setting
@@
Theorem A.2.
Fix $d>0$. Consider any PE problem $(\bar{s},M,\mathcal{M},\pi_{M},\Pi)$ satisfying Assumptions 4.1 and A.1 for the same feature map $\phi$, then there exists a fully-adaptive learner that solves the PE problem exactly in at most $d$ queries.
@@
Theorem A.2 together with Theorem A.3 shows that exactly $d$ queries are optimal for solving a PE problem under our feedback model and Assumption A.1. The lower-bound is to be expected because the learner is operating in a $d$-dimensional feature space so has to learn in $d$ directions. However, it is interesting that the structure imposed by Assumption A.1 on the learner’s capacity to explore the feature space is sufficient for the learner to fully solve the problem in only $d$ queries. A similar assumption was studied by Jia et al. (2023) to obtain a sample-efficient algorithm for BPI in finite-horizon MDPs. We can obtain this result for PE with a simple analysis because the learner can exploit that the action-value of a policy is the fixed point of a linear operator, the Bellman evaluation operator. An equivalent approach does not work for BPI since the action-value of the optimal policy is not the fixed point of a linear operator.
The proofs of the theorems in this section can be found in Appendix F.
Appendix B Hyper-spherical caps and sectors for subspaces
@@
Beyond the extension to subspaces of multiple dimensions, Definition B.2 differs from Definition B.1 in two ways. It is a hyper-spherical sector rather than cap, which does not restrict the vectors to be close to the boundary of $\mathcal{B}$. It is two sided meaning that if $x\in\overset{\triangle}{\mathcal{C}}_{\gamma}(H)$, then $-v\in\overset{\triangle}{\mathcal{C}}_{\gamma}(H)$. Note that the subspace $H$ is defined on $\mathbb{R}^{d}$ but $\overset{\triangle}{\mathcal{C}}_{\gamma}(H)\subset\mathcal{B}$.
@@
Appendix C Subspace Packing
C.1 Preliminaries
@@
C.2 A subspace packing bound
Lemma C.2 (Soleymani & Mahdavifar (2021), Theorem 4).
@@ C.2 A subspace packing bound
C.3 Existence of an isolated subspace
The following lemma is the key result for the construction of the class of MDPs used in the proof of our main results. It establishes that if a number of points $n$ is “small”, then for any $n$ points there exists a subspace whose $\gamma$-sector contains none of the $n$ points. The environment can erase information along this subspace (see Section 5). See Appendix B for the definition of $\overset{\triangle}{\mathcal{C}}_{\gamma}(H)$ for a subspace $H$. The proof is given in Appendix C.3.2.
Lemma C.3.
Fix $d=2^{N}$ for some $N\geq 8$. Consider $D=\{y_{1},...,y_{n}\}$, a set of $n$ points s.t $y_{i}\in\mathcal{B}$ for all $i\in[n]$. If
@@ C.3 Existence of an isolated subspace
then there exists a subspace $A\in\mathcal{G}_{2^{\lceil N/4\rceil},d}(\mathbb{R})$ of dimension $2^{\lceil N/4\rceil}$ s.t
@@
C.3.1 Preliminary Lemmas
The proof of Lemma C.3 relies on the following lemmas. See Appendix C for the definition of $\mathcal{G}_{m,d}(\mathbb{R})$ and the definition of a subspace packing $\mathcal{C}$.
Lemma C.4.
Fix $d=2^{N}$ for some $N\in\mathbb{N}$. If there exists a packing $\mathcal{C}$ in $G_{m,d}(\mathbb{R})$ of size $|\mathcal{C}|\geq n+1$ s.t $d_{\min}(\mathcal{C})\geq\sqrt{m-g(\gamma)^{2}}$, then given a set $\{y_{1},...,y_{n}\}$ of $n$ queries s.t $y_{i}\in\mathcal{B}$ for all $i\in[n]$, there exists a subspace $H\in G_{m,d}(\mathbb{R})$ of dimension $m$ s.t $y_{i}\notin\overset{\triangle}{\mathcal{C}}_{\gamma}(A)$ for all $i$.
This lemma establishes that if $n+1$ subspaces are sufficiently far in terms of chordal distance, then for any $n$ points there is a subspace whose $\gamma$-sector does not contain any of the $n$ points. The proof is given in Appendix C.3.3
Lemma C.5.
Consider $(n+1)$ subspaces $A_{1},...,A_{n+1}$ of dimension $m$ s.t for any $i\neq j$,
Given $n$ vectors $y_{1},...,y_{n}\in\mathcal{B}$ (w.l.o.g. all unit norm), then there is a subspace $H\in\{A_{1},...,A_{n+1}\}$ s.t for all $y\in\{y_{1},...,y_{n}\}$,

### v2

@@ 2 Preliminaries
A Markov Decision Process (MDP) (Puterman, 1994) is a discrete-time stochastic process comprised of a set of states $\mathcal{S}$, a set of actions $\mathcal{A}=\bigcup_{s\in\mathcal{S}}\{\mathcal{A}_{s}\}$ where $\mathcal{A}_{s}$ is the action space in state $s\in\mathcal{S}$ and, for each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}_{s}$, a next-state transition function given by a measure $p(\cdot|s,a)$ and a (deterministic) reward function $r(s,a)\in[-1,1]$ (in fact, even the value functions defined below are in $[-1,1]$ for our constructions). In a state $s$, an agent chooses an action $a$, receives a reward $r(s,a)$ and transitions to a new state according to $p(\cdot|s,a)$. Once in the new state, the process continues. The actions chosen by an agent are formalised by policies. A deterministic policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ is a mapping from a state to an action. In each state $s\in\mathcal{S}$, an agent following policy $\pi$ chooses action $\pi(s)\in\mathcal{A}_{s}$. We do not consider stochastic policies.
@@ 4 Main Results
In this section, we present our main results: lower-bounds on the number of rounds $K$ for sample-efficient algorithms. First, we state some assumptions. We assume $\gamma>\sqrt{3/4}$. We also consider a linear representation of the action-values for a known feature map of state-action pairs. This is a form of linear function approximation that is strictly more general than linear MDPs (Zanette et al., 2020), which assume the reward and transition functions are linearly representable. Specifically:
@@
This first assumption is used for PE problems. The second is stronger as it concerns the action-value of every policy (not just policies in $\Pi$) and in particular it holds for $Q^{\star}$. We assume the learner is aware when these assumptions hold. Finally, we formally define a sample-efficient learner:
@@ 4.1 Policy-Induced Queries
We first present a result for PE under policy-induced queries. Since all MDPs in the class used in the proof share the same dynamics, policy-induced queries are equivalent to policy-free queries (see Remark 3.6) and the result holds for both. The full proof can be found in Appendix E.
@@
In the class of MDPs used for Theorem 4.4 we can hide information about $M\in\mathcal{M}$ in the target-policy for PE but cannot for BPI. Instead, we could hide information in the transitions but this can be revealed by following policy trajectories (policy-induced queries) in our constructions. In the next section, we restrict the learner to policy-free queries and provide results for both PE and BPI. Note that a $(1,1/2)$-sound learner performs poorly since the MDP class has value functions in $[-1,1]$.
@@ 4.2 Policy-Free Queries
We now consider only policy-free queries, which gives the environment freedom to hide information in the transition function of the MDP and leads to lower-bounds for PE and BPI that hold for the stronger Assumption 4.2 of all-policy realizability. The full proof can be found in Appendix F.
@@ 4.3 Discussion
Our lower-bounds no longer hold if the action space $\mathcal{A}$ is finite or with coverage assumptions on the collected data (for BPI). While it may be necessary to have coverage assumptions in an offline setting, this assumption need not be made when data is collected adaptively, as in our setting.
Furthermore, in Appendix B we provide results for the fully adaptive setting where the number of batches $K$ is equal to the number of total queries $n$ (i.e. each batch contains a single query). We show that if the feature-space covers $\mathcal{B}$, there is a learner that solves any realizable (Assumption 4.1) PE problem in $d$ queries. Assuming a known target policy $\pi$, a linear dependence on $d$ was already known to be possible using roll-outs from $\pi$ (Lattimore et al., 2020), however the dependence on $d$ was coupled with other quantities such as the effective horizon $1/(1-\gamma)$ and desired accuracy $\varepsilon$. Our result states that $d$-queries are sufficient to find $Q^{\pi}$ exactly ($\varepsilon=0$), independently of $\gamma$. Our result relies heavily on the condition that the learner observes the transition function $p_{M}(\cdot|s,a)$ rather than a sample $s^{\prime}\sim p_{M}(\cdot|s,a)$ for a query $(s,a)$ (see Section 3.3), though our result under this condition is strong since using this condition with the roll-outs from $\pi$ would not give exact convergence in $d$-queries. We also provide a matching lower-bound (that also holds for BPI). These results serve to illustrate the trade-off between sample-efficiency and low-adaptivity for PE under our framework. Full-adaptivity allows low sample-complexity, while reducing adaptivity below $K=o(\log\log d)$ comes at the cost of high sample-complexity (losing sample-efficiency).
4.4 Related Works
Please refer to Appendix A for a discussion of works related to low-switching cost (Qiao et al., 2022; Bai et al., 2019; Zhang et al., 2020; Gao et al., 2021; Wang et al., 2021; Qiao & Wang, 2023; Qiao et al., 2023) and policy fine-tuning (Xie et al., 2021; Zhang & Zanette, 2023).
@@ 5 Proof Sketch
In this section, we provide intuition for the proof of Theorem 4.4. We extend the ideas of Zanette (2021) beyond offline RL to our multi-batch problem (see the end of this section for a comparison). We consider the PE problem with policy-free queries under Assumption 4.1 for feature vectors $\phi(\cdot,\cdot)$ covering the unit Euclidean ball $\mathcal{B}$ (see Figure 1). The intuition for Theorem 4.5 is closely related.
@@
where $\mathcal{C}_{\gamma}(w)=\{x\in\mathcal{B}:x^{T}w/\|w\|_{2}>\gamma\}$ is the $\gamma$-hyperspherical cap of $w$. For a $w\in\partial\mathcal{B}$ and a carefully designed target policy $\pi_{w}$, we show that Assumption 4.1 holds. We also show with the reasoning described above that if $n_{k}=poly(d)$ for all $k$ and $K=o(\log\log d)$, then none of the learner’s queries are in $\mathcal{C}_{\gamma}(w)\cup\mathcal{C}_{\gamma}(-w)$. Therefore the feedback observed contains no information about the sign of the rewards in $\mathcal{C}_{\gamma}(w)\cup\mathcal{C}_{\gamma}(-w)$. $M_{w,+}$ is indistinguishable from $M_{w,-}$. Since $Q_{M_{w,+}}^{\pi_{w}}(\bar{s},w)=1$ and $Q_{M_{w,-}}^{\pi_{w}}(\bar{s},w)=-1$, the learner must incur an error of $1$ with probability at least $1/2$. All the details are in Appendix E, including further illustrations in Appendix E.5. The construction for Theorem 4.5 is similar but the transitions differ across MDPs (see Appendix F).
To improve the lower-bound with the current construction, we require the existence of a subspace packing (of size exponential in $d$ and subspace dimension $cd$ (instead of $d^{1/4}$), for some $c<1$) with minimal chordal distance (see Appendix D) between subspaces $d_{\min}(\mathcal{C})\geq\sqrt{d-(2\gamma^{2}-1)^{2}}$. After $k$ rounds, information would be missing along a subspace of dimension $c^{k}d$, (instead of $d^{1/4^{k}}$) from which we could get a $\log d$ lower-bound. As far as we are aware, such a result is not available in the literature, nor is a subspace covering result that would rule out the possibility of such a packing. We highlight that our MDP construction can be combined with any subspace packing result, paving the way for improved lower-bounds should new subspace packing procedures be developed.
Comparison to Zanette (2021): The work of Zanette (2021) erases information along a 1-dimensional subspace ($X$ is of rank $d-1$). Therefore after having observed one round of feedback, a single additional query is sufficient to distinguish the MDP and solve the problem. Our constructions erase information along $m$-dimensional subspaces where we attempt to maximise $m$ so that even after having observed feedback revealing this subspace, a polynomial number of queries remains insufficient to distinguish the MDP, more adaptive rounds are needed.
@@ 6 Conclusion
It remains unclear if the $\log\log d$ dependence on $d$ is tight. An upper-bound similar to the one we have given for the fully-adaptive PE problem in Appendix B could be established by developing new tools in the theory of subspace covering. These would formalise the number of directions for which the learner can prevent information erasure. It is also unclear if Theorem 4.4 under policy-induced queries also holds for BPI or if the sample-efficiency of BPI-algorithms in low-adaptivity settings differs for policy-induced and policy-free queries. We leave these as future work.
Acknowledgments and Disclosure of Funding
EJ is funded by EPSRC through the Modern Statistics and Statistical Machine Learning (StatML) CDT (grant no. EP/S023151/1) and thanks G-Research for partly funding attendance to ICLR 2024.
We would like to thank the reviewers and meta-reviewers for their time and feedback, which led to a better version of this paper. We would also like to thank Nathan Mankovich and Hugo Chu for helpful discussions.
Appendix A Further Related Works
Appendix B Bounds for the Fully Adaptive Setting
@@
Theorem B.2.
Fix $d>0$. Consider any PE problem $(\bar{s},M,\mathcal{M},\pi_{M},\Pi)$ satisfying Assumptions 4.1 and B.1 for the same feature map $\phi$, then there exists a fully-adaptive learner that solves the PE problem exactly in at most $d$ queries.
@@
Theorem B.2 together with Theorem B.3 shows that exactly $d$ queries are optimal for solving a PE problem under our feedback model and Assumption B.1. The lower-bound is to be expected because the learner is operating in a $d$-dimensional feature space so has to learn in $d$ directions. However, it is interesting that the structure imposed by Assumption B.1 on the learner’s capacity to explore the feature space is sufficient for the learner to fully solve the problem in only $d$ queries. A similar assumption was studied by Jia et al. (2023) to obtain a sample-efficient algorithm for BPI in finite-horizon MDPs. We can obtain this result for PE with a simple analysis because the learner can exploit that the action-value of a policy is the fixed point of a linear operator, the Bellman evaluation operator. An equivalent approach does not work for BPI since the action-value of the optimal policy is not the fixed point of a linear operator.
The proofs of the theorems in this section can be found in Appendix G.
Appendix C Hyper-spherical caps and sectors for subspaces
@@
See Figure 2 for an illustration.
Figure 2: Illustration of a hyperspherical sector of a $1$-dimensional subspace (left) and a $2$-dimensional subspace (right - all vectors whose direction is within the two pink bands). In both cases, the subspace $H$ is in yellow.
Beyond the extension to subspaces of multiple dimensions, Definition C.2 differs from Definition C.1 in two ways. It is a hyper-spherical sector rather than cap, which does not restrict the vectors to be close to the boundary of $\mathcal{B}$. It is two sided meaning that if $x\in\overset{\triangle}{\mathcal{C}}_{\gamma}(H)$, then $-v\in\overset{\triangle}{\mathcal{C}}_{\gamma}(H)$. Note that the subspace $H$ is defined on $\mathbb{R}^{d}$ but $\overset{\triangle}{\mathcal{C}}_{\gamma}(H)\subset\mathcal{B}$.
@@
We also note that a vector $x$ is not in $\overset{\triangle}{\mathcal{C}}_{\gamma}(H)$ if the subspace $H$ does not intersect $\overset{\triangle}{\mathcal{C}}_{\gamma}(\langle x\rangle)$, the $\gamma$-hyperspherical sector of $\langle x\rangle$. This allows us to work with the $\gamma$-hyperspherical sectors of one dimensional subspaces (span of a single vector) instead of subspaces of multiple dimensions. We will also abuse notation and write $\overset{\triangle}{\mathcal{C}}_{\gamma}(x)$ instead of $\overset{\triangle}{\mathcal{C}}_{\gamma}(\langle x\rangle)$
Appendix D Subspace Packing
D.1 Preliminaries
@@
D.2 A subspace packing bound
Lemma D.2 (Soleymani & Mahdavifar (2021), Theorem 4).
@@ D.2 A subspace packing bound
D.3 Existence of an isolated subspace
The following lemma is the key result for the construction of the class of MDPs used in the proof of our main results. It establishes that if a number of points $n$ is “small”, then for any $n$ points there exists a subspace whose $\gamma$-sector contains none of the $n$ points. The environment can erase information along this subspace (see Section 5). See Appendix C for the definition of $\overset{\triangle}{\mathcal{C}}_{\gamma}(H)$ for a subspace $H$. The proof is given in Appendix D.3.2.
Throughout, we will use $g(\gamma)=2\gamma^{2}-1$.
Lemma D.3.
Fix $d=2^{N}$ for some $N\geq 8$. Consider $D=\{y_{1},...,y_{n}\}$, a set of $n$ points s.t. $y_{i}\in\mathcal{B}$ for all $i\in[n]$. If
@@ D.3 Existence of an isolated subspace
then there exists a subspace $A\in\mathcal{G}_{2^{\lceil N/4\rceil},d}(\mathbb{R})$ of dimension $2^{\lceil N/4\rceil}$ s.t.
@@
D.3.1 Preliminary Lemmas
The proof of Lemma D.3 relies on the following lemmas. See Appendix D for the definition of $\mathcal{G}_{m,d}(\mathbb{R})$ and the definition of a subspace packing $\mathcal{C}$.
Lemma D.4.
Fix $d=2^{N}$ for some $N\in\mathbb{N}$. If there exists a packing $\mathcal{C}$ in $G_{m,d}(\mathbb{R})$ of size $|\mathcal{C}|\geq n+1$ s.t. $d_{\min}(\mathcal{C})\geq\sqrt{m-g(\gamma)^{2}}$, then given a set $\{y_{1},...,y_{n}\}$ of $n$ queries s.t. $y_{i}\in\mathcal{B}$ for all $i\in[n]$, there exists a subspace $H\in G_{m,d}(\mathbb{R})$ of dimension $m$ s.t. $y_{i}\notin\overset{\triangle}{\mathcal{C}}_{\gamma}(A)$ for all $i$.
This lemma establishes that if $n+1$ subspaces are sufficiently far in terms of chordal distance, then for any $n$ points there is a subspace whose $\gamma$-sector does not contain any of the $n$ points. The proof is given in Appendix D.3.3
Lemma D.5.
Consider $(n+1)$ subspaces $A_{1},...,A_{n+1}$ of dimension $m$ s.t. for any $i\neq j$,
$\max_{x\in A_{i},z\in A_{j}}\frac{x^{T}z}{\|x\|_{2}\|z\|_{2}}<g(\gamma).$
Given $n$ vectors $y_{1},...,y_{n}\in\mathcal{B}$ (w.l.o.g. all unit norm), then there is a subspace $H\in\{A_{1},...,A_{n+1}\}$ s.t. for all $y\in\{y_{1},...,y_{n}\}$,